OpenStudy (anonymous):

4. Jose received $30 for his birthday that he put in his piggy bank. Each week, he puts 4 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks. His sister Marisol makes a table of reasonable values for w to determine how much money Andre will eventually have. If one of the values for w Marisol selects is 36, what would be her corresponding entry for D? (Points : 1) 466 174 70 39

OpenStudy (campbell_st):

\[D = 4\times36 + 30\] just calculate it

OpenStudy (sleepyjess):

Remember your order of operations PEMDAS (Parenthesis, Exponent, Multiplication, Division, Addition, Subtraction D = 4 x 36 + 30 So Multipy first D = (4 x 36) + 30 D = 144 + 30